OpenStudy (anonymous):

3(2y - 8) - 2y(5 - y) -2y2+ 16y - 24 -5y - 8 2y2- 4y - 24 2y2+ 16y - 24

OpenStudy (anonymous):

help sos]

OpenStudy (whpalmer4):

Just use the distributive property over and over until no parentheses remain, then collect like terms to simplify. \[3(2y-8)-2y(5-y) = 3*2y-3*8-2y(5-y) = 6y-24-2y(5-y)\]I'll let you do the rest. Don't forget that if you have negative number on the outside of the parentheses, it gets applied to each term within...

OpenStudy (anonymous):

what!!!!

OpenStudy (whpalmer4):

distributive property: \[a(b+c) = ab + ac\]

OpenStudy (anonymous):

ur not getting the medal ur not helping

OpenStudy (whpalmer4):

I'm not giving you the answer, if that's what you want. Medal, schmedal.

OpenStudy (uri):

He pretty much helped you @wyattp96,And one more thing is this your second account?

OpenStudy (anonymous):

no

OpenStudy (whpalmer4):

What don't you understand about how to do this?

OpenStudy (anonymous):

its hard

OpenStudy (uri):

I think he doesn't understand how you have written it. @whpalmer4 maybe write in verticle.

OpenStudy (anonymous):

plz

OpenStudy (whpalmer4):

Do you understand the distributive property of multiplication? \[a(b+c) = a*b + a*c\]Think of it with numbers: 10*16 = 160 10*(4+12) = 160 10*4 + 10*12 = 40 + 120 = 160

OpenStudy (whpalmer4):

For the first part of your equation, you have \[3(2y - 8)\]We multiply 3 and 2y, which gives 6y. We multiply 3 and -8, which gives -24. We add them together, and get \(6y-24\) so \[3(2y-8) = 6y-24\] Yes?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

screw it im guessing

OpenStudy (whpalmer4):

If you don't understand this, math classes are going to be a torment...good luck!

OpenStudy (anonymous):

...

OpenStudy (uri):

3(2y - 8) - 2y(5 - y) = 6y - 24 - 10y + 2y^2 Understand?

OpenStudy (uri):

Put y^2 values together y values together and -24 seperate..

OpenStudy (uri):

@wyattp17